THE FANTASIO SET-UP (II): HIGH RESOLUTION OVERTONE SPECTROSCOPY OF ACETYLENE CONTAINING VAN DER WAALS DIMERS

C. LAUZIN, J. DEMAISON, K. DIDRICHE, P. MACKO, J. LIEVIN, \underline{M. HERMAN}, Service de Chimie quantique et Photophysique CP160/09, Faculté des Sciences, Université Libre de Bruxelles (U.L.B.), Av. Roosevelt, 50, B-1050, Bruxelles, Belgium; A. PERRIN, Laboratoire Interuniversitaire des Systèmes Atmosphériques (LISA) CNRS-UMR 7583 Université Paris Est et Paris 7, Faculté des Sciences et Technologie, 61, avenue du Général de Gaulle 94010 Créteil Cédex, France; W.J. LAF-FERTY, Optical Technology Division, National Institute for Standards and Technology, Gaithersburg, MD 20899-8441, USA.

We have used the CW-CRDS facility in the FANTASIO set-up described in the previous presentation to investigate acetylene-containing dimers formed in a supersonic expansion, in the 1.5 microns range. Ab initio calculations were also performed in some cases. The acetylene partners so far include argon^{a}, carbon dioxide^{b} and nitrous oxide^{c}. The results will be illustrated.

aC. Lauzin, K. Didriche, P. Macko, J. Demaison, J. Liévin, and M. Herman, J. Phys. Chem. A, in press (2009).

bC. Lauzin, K. Didriche, J. Liévin, M. Herman, and A. Perrin, J. Chem. Phys, submitted (2009).

cK. Didriche, C. Lauzin, P. Macko, M. Herman, and W.J. Lafferty, Chem. Phys. Letters, doi:10.1016/j.cplett.2008.12.037 (2008).